OpenStudy (anonymous):

Solve for b: 2b/4 + b/3 =10

OpenStudy (anonymous):

B=12 :)

OpenStudy (anonymous):

here it goes: 4b/2 + 3 =11+b 4b/2= 11 + b - 3 (transposing 3 on right side ) 2b = 8 + b ( since 4b/2 = 2b , i hope u know how to divide , 2 divides 4 of the numerator ) 2b - b = 8 (transposing b on left side) b=8 which is the required answer.

OpenStudy (emmamink):

b equals 12. Plug it into the equation to see for yourself.

OpenStudy (emmamink):

Hope that helped!

OpenStudy (gschibby):

Wrong approach...

OpenStudy (anonymous):

Thank you @Rawr_My_Love143 <3

OpenStudy (anonymous):

I got 12 (: Thank you

OpenStudy (gschibby):

\[\frac{ 2b }{ 4 }+\frac{ b }{ 3 }=10\]Multiply everything with the common denominator, in this case 3*4=12:\[(3*2b)+4b=120\]Add all the b's:\[6b+4b=10b=120\]Now divide with with 10 on both sides:\[b=12\]And there you have it! ;)

OpenStudy (like1234):

Solution for 2b/4+b/3=10 equation: b in (-oo:+oo) (2*b)/4+b/3 = 10 // - 10 (2*b)/4+b/3-10 = 0 5/6*b-10 = 0 // + 10 5/6*b = 10 // : 5/6 b = 10/5/6 b = 12 b = 12 http://www.geteasysolution.com/2b/4+b/3=10